# Measurement of the scintillation resolution in liquid xenon and its impact for future segmented calorimeters

### v5

@@ Acknowledgements.
- [1] NaI(Tl) properties, Online, accessed: 2023-03-30.
- [2] CsI(Tl) properties, Online, accessed: 2023-03-30.
- [3] LYSO(Ce) properties, Online, accessed: 2023-03-30.
- [4] BGO properties, Online, accessed: 2023-03-30.
- Yeh and Zhu [2022] M. Yeh and R.-Y. Zhu, Materials for future calorimeters (2022).
- Vandeberghe et al. [2020] S. Vandeberghe, P. Moskal, and J. Karp, State of the art in total body pet, EJNMMI Physics 7, 35 (2020).
- Westerwoudt et al. [2014] V. Westerwoudt, M. Conti, and L. Eriksson, Advantages of improved time resolution for TOF PET at very low statistics, IEEE Trans. Nucl. Sci. 61, 126 (2014).
- Conti and Bendriem [2019] M. Conti and B. Bendriem, The new opportunities for high time resolution clinical TOF PET, Clinical and Translational Imaging 7, 139 (2019).
- Cherry et al. [2018] S. R. Cherry et al., Total-body PET: Maximizing sensitivity to create new opportunities for clinical research and patient care, J. Nucl. Med. 59, 3 (2018).
- Badawi et al. [2019] R. D. Badawi et al., First human imaging studies with the EXPLORER total-body PET scanner, J. Nucl. Med. 60, 299 (2019).
- Aprile and Doke [2010] E. Aprile and T. Doke, Liquid xenon detectors for particle physics and astrophysics, Rev. Mod. Phys 82, 2053 (2010), arXiv:1207.2292 [physics.ins-det] .
- Fujii et al. [2015] K. Fujii et al., High-accuracy measurement of the emission spectrum of liquid xenon in the vacuum ultraviolet region, Nuclear Instruments and Methods in Physics Research Section A: Accelerators, Spectrometers, Detectors and Associated Equipment 795, 293 (2015).
- Chepel and Araujo [2013] V. Chepel and H. Araujo, Liquid noble gas detectors for low energy particle physics, JINST 8, R04001, arXiv:1207.2292 [physics.ins-det] .
- Kubota et al. [1978] S. Kubota, A. Nakamoto, T. Takahashi, T. Hamada, E. Shibamura, M. Miyajima, K. Masuda, and T. Doke, Recombination luminescence in liquid argon and in liquid xenon, Phys. Rev. B 17, 2762 (1978).
- Hogenbirk et al. [2018] E. Hogenbirk et al., Field dependence of electronic recoil signals in a dual-phase liquid xenon time projection chamber, Journal of Instrumentation 13 (10), P10031–P10031.
- Gallucci [2009] G. Gallucci, The MEG liquid xenon calorimeter, Journal of Physics: Conference Series 160, 012011 (2009).
- Anton et al. [2019] G. Anton et al. (EXO-200 Collaboration), Search for neutrinoless double- decay with the complete EXO-200 dataset, Phys. Rev. Lett. 123, 161802 (2019).
- Aprile et al. [2020] E. Aprile et al., Energy resolution and linearity of XENON1T in the mev energy range, Eur. Phys. J. C 80, 785 (2020).
- Adhikari et al. [2021] G. Adhikari et al., Journal of Physics G: Nuclear and Particle Physics 49, 015104 (2021).
@@
- Aalbers et al. [2016] J. Aalbers et al., DARWIN: towards the ultimate dark matter detector, Journal of Cosmology and Astroparticle Physics 2016 (11), 017.
- Lavoie [1976] L. Lavoie, Liquid xenon scintillators for imaging of positron emitters, Med. Phys. 3 (5), 283 (1976).
- Chepel et al. [1999] V. Chepel et al., The liquid xenon detector for PET: recent results, IEEE Transactions on Nuclear Science NS-46, 1038 (1999).
- Crespo et al. [1999] P. Crespo et al., Pulse processing for the PET liquid xenon Multiwire Ionisation Chamber, IEEE Transactions on Nuclear Science 47, 2119 (1999).
- Doke et al. [2006] T. Doke, J. Kikuchi, and F. Nishikido, Time-of-flight positron emission tomography using liquid xenon scintillation, Nucl. Instrum. Methods A 569, 863 (2006).
- Miceli [2011] A. Miceli, Liquid xenon detectors for positron emission tomography, J. Phys. Conf. Ser. 312, 062006 (2011).
- Gallego Manzano et al. [2015] L. Gallego Manzano et al., XEMIS: A liquid xenon detector for medical imaging, Nucl. Instrum. Meth. A787, 89 (2015).
- Chepel [2007] V. Chepel, Liquid xenon detectors for medical imaging, Revista do Detua 4 (2007).
- Gomez-Cadenas et al. [2016] J. J. Gomez-Cadenas et al., Investigation of the Coincidence Resolving Time performance of a PET scanner based on liquid xenon: A Monte Carlo study, JINST 11 (09), P09011, arXiv:1604.04106 [physics.ins-det] .
- Gomez-Cadenas et al. [2017] J. J. Gomez-Cadenas, J. M. Benlloch-Rodríguez, and P. Ferrario, Monte Carlo study of the Coincidence Resolving Time of a liquid xenon PET scanner, using Cherenkov radiation, JINST 12 (08), P08023, arXiv:1706.07629 [physics.ins-det] .
- Herrero-Bosch et al. [2018] V. Herrero-Bosch et al., PETALO read-out: A novel approach for data acquisition systems in PET applications, in 2018 IEEE Nuclear Science Symposium and Medical Imaging Conference (NSS/MIC) (2018).
- Renner et al. [2019] J. Renner et al., Processing of Compton events in the PETALO readout system, in 2019 IEEE Nuclear Science Symposium (NSS) and Medical Imaging Conference (MIC) (2019) pp. 1–7, arXiv:2001.04724 [physics.ins-det] .
- Ferrario et al. [2018] P. Ferrario et al., PETALO: Time-of-Flight PET with liquid xenon, in 2018 IEEE Nuclear Science Symposium and Medical Imaging Conference (NSS/MIC) (2018) arXiv:1911.10994 [physics.ins-det] .
- Renner et al. [2022] J. Renner et al., Monte carlo characterization of PETALO, a full-body liquid xenon-based PET detector, Journal of Instrumentation 17, P05044.
- Salor-Iguiñiz et al. [2025] N. Salor-Iguiñiz et al., Evaluation of coincidence time resolution in a liquid xenon detector with silicon photomultipliers, Phys. Rev. Res. 7, 033089 (2025).
- Doke et al. [2003] T. Doke, R. Sawada, and H. Tawara, Non-proportionality of the scintillation yield in liquid xenon and its effect on the energy resolution for gamma-rays, Technique and Application of Xenon Detectors , 17 (2003).
- Szydagis et al. [2018] M. Szydagis et al., Noble element simulation technique v2.0 (version v2.0.0), Zenodo 10.5281/zenodo.1314669 (2018).
- Yamashita et al. [2004] M. Yamashita et al., Scintillation response of liquid xe surrounded by ptfe reflector for gamma rays, Nuclear Instruments and Methods in Physics Research Section A: Accelerators, Spectrometers, Detectors and Associated Equipment 535, 692 (2004).
- Plante et al. [2011] G. Plante, E. Aprile, R. Budnik, B. Choi, K.-L. Giboni, L. W. Goetzke, R. F. Lang, K. E. Lim, and A. J. Melgarejo Fernandez, New measurement of the scintillation efficiency of low-energy nuclear recoils in liquid xenon, Phys. Rev. C 84, 045805 (2011).
- Ni et al. [2006] K. Ni et al., Gamma Ray Spectroscopy with Scintillation Light in Liquid Xenon, JINST 1, P09004, arXiv:physics/0608034 [physics] .
- Baldini et al. [2005] A. Baldini et al., Absorption of scintillation light in a 100 l liquid xenon gamma-ray detector and expected detector performance, Nuclear Instruments and Methods in Physics Research Section A 545, 753 (2005).
- Akerib et al. [2013] D. Akerib et al., Technical results from the surface run of the lux dark matter experiment, Astroparticle Physics 45, 34 (2013).
- Di Francesco et al. [2016] A. Di Francesco et al., Tofpet2: a high- performance asic for time and amplitude measurements of sipm signals in time-of-flight applications, JINST 11 (3), C03042.
@@
- Agostinelli et al. [2003] S. Agostinelli et al., GEANT4 - a simulation toolkit, Nucl. Instrum. Meth. A 506, 250 (2003).
@@
- Baldini et al. [2006] A. Baldini et al., Liquid xenon scintillation calorimetry and xe optical properties, IEEE Transactions on Dielectrics and Electrical Insulation 13, 547 (2006).
- Solovov et al. [2004] V. Solovov et al., Measurement of the refractive index and attenuation length of liquid xenon for its scintillation light, Nucl. Instrum. Meth. A 516, 462 (2004).
@@
- [46] Linearity curve for Hamamatsu SiPMs, Online, accessed: 2023-07-18.

### v6

@@ Acknowledgements.
- [1] (2016) DARWIN: towards the ultimate dark matter detector. Journal of Cosmology and Astroparticle Physics 2016 (11), pp. 017–017. Cited by: §I.
- [2] (2021) nEXO: neutrinoless double beta decay search beyond 1028 year half-life sensitivity. Journal of Physics G: Nuclear and Particle Physics 49 (1), pp. 015104. Cited by: §I.
@@
- [3] (2003) GEANT4 - a simulation toolkit. Nucl. Instrum. Meth. A 506, pp. 250. Cited by: Appendix A, §III.
- [4] (2018) Calibration, event reconstruction, data analysis, and limit calculation for the LUX dark matter experiment. Phys. Rev. D 97, pp. 102008. Cited by: Appendix A.
- [5] (2013) Technical results from the surface run of the lux dark matter experiment. Astroparticle Physics 45, pp. 34–43. Cited by: §II.
- [6] (2019) Search for neutrinoless double- decay with the complete EXO-200 dataset. Phys. Rev. Lett. 123, pp. 161802. Cited by: §I.
- [7] (2010) Liquid xenon detectors for particle physics and astrophysics. Rev. Mod. Phys 82, pp. 2053. Cited by: §I, §I.
- [8] (2020) Energy resolution and linearity of XENON1T in the mev energy range. Eur. Phys. J. C 80, pp. 785. Cited by: §I.
- [9] (2019) First human imaging studies with the EXPLORER total-body PET scanner. J. Nucl. Med. 60, pp. 299–303. Cited by: §I.
- [10] (2005) Absorption of scintillation light in a 100 l liquid xenon gamma-ray detector and expected detector performance. Nuclear Instruments and Methods in Physics Research Section A 545, pp. 753–764. Cited by: Appendix A, §II.
@@
- [11] (2006) Liquid xenon scintillation calorimetry and xe optical properties. IEEE Transactions on Dielectrics and Electrical Insulation 13 (3), pp. 547–555. Cited by: Appendix A, §III.
- [12] BGO properties. Note: Online, accessed: 2023-03-30 External Links: Link Cited by: §I.
- [13] (2013) Liquid noble gas detectors for low energy particle physics. JINST 8, pp. R04001. Cited by: §I.
- [14] (1999) The liquid xenon detector for PET: recent results. IEEE Transactions on Nuclear Science NS-46, pp. 1038–1044. Cited by: §I.
- [15] (2007) Liquid xenon detectors for medical imaging. Revista do Detua 4 (7). Cited by: §I.
- [16] (2018) Total-body PET: maximizing sensitivity to create new opportunities for clinical research and patient care. J. Nucl. Med. 59 (1), pp. 3–12. Cited by: §I.
- [17] (2019) The new opportunities for high time resolution clinical TOF PET. Clinical and Translational Imaging 7, pp. 139 – 147. Cited by: §I.
- [18] (1999) Pulse processing for the PET liquid xenon Multiwire Ionisation Chamber. IEEE Transactions on Nuclear Science 47, pp. 2119–2126. Cited by: §I.
- [19] CsI(Tl) properties. Note: Online, accessed: 2023-03-30 External Links: Link Cited by: §I.
- [20] (2016) Tofpet2: a high- performance asic for time and amplitude measurements of sipm signals in time-of-flight applications. JINST 11 (3), pp. C03042. Cited by: §II.
- [21] (2006) Time-of-flight positron emission tomography using liquid xenon scintillation. Nucl. Instrum. Methods A 569, pp. 863. Cited by: §I.
- [22] (2003) NON-proportionality of the scintillation yield in liquid xenon and its effect on the energy resolution for gamma-rays. Technique and Application of Xenon Detectors, pp. 17–27. Cited by: Figure 2, §I, §V, §V, §VI.
- [23] (2018-11) PETALO: Time-of-Flight PET with liquid xenon. In 2018 IEEE Nuclear Science Symposium and Medical Imaging Conference (NSS/MIC), Cited by: §I.
- [24] (2015) High-accuracy measurement of the emission spectrum of liquid xenon in the vacuum ultraviolet region. Nuclear Instruments and Methods in Physics Research Section A: Accelerators, Spectrometers, Detectors and Associated Equipment 795, pp. 293–297. Cited by: §I.
- [25] (2015) XEMIS: a liquid xenon detector for medical imaging. Nucl. Instrum. Meth. A787, pp. 89–93. Cited by: §I.
- [26] (2009) The MEG liquid xenon calorimeter. Journal of Physics: Conference Series 160 (1), pp. 012011. Cited by: §I.
- [27] (2016) Investigation of the Coincidence Resolving Time performance of a PET scanner based on liquid xenon: A Monte Carlo study. JINST 11 (09), pp. P09011. Cited by: §I.
- [28] (2017) Monte Carlo study of the Coincidence Resolving Time of a liquid xenon PET scanner, using Cherenkov radiation. JINST 12 (08), pp. P08023. Cited by: §I.
- [29] (2018) PETALO read-out: A novel approach for data acquisition systems in PET applications. In 2018 IEEE Nuclear Science Symposium and Medical Imaging Conference (NSS/MIC), Cited by: §I, §II.
- [30] (2018) Field dependence of electronic recoil signals in a dual-phase liquid xenon time projection chamber. Journal of Instrumentation 13 (10), pp. P10031–P10031. Cited by: §I.
- [31] (2020) Measurements of angle-resolved reflectivity of PTFE in liquid xenon with IBEX. The European Physical Journal C 80 (3), pp. 262. Cited by: Appendix A.
- [32] (1978) Recombination luminescence in liquid argon and in liquid xenon. Phys. Rev. B 17, pp. 2762–2765. Cited by: §I.
- [33] (1976) Liquid xenon scintillators for imaging of positron emitters. Med. Phys. 3 (5), pp. 283–293. Cited by: §I.
@@
- [34] Linearity curve for Hamamatsu SiPMs. Note: Online, accessed: 2023-07-18 External Links: Link Cited by: §III.
- [35] LYSO(Ce) properties. Note: Online, accessed: 2023-03-30 External Links: Link Cited by: §I.
- [36] (2011) Liquid xenon detectors for positron emission tomography. J. Phys. Conf. Ser. 312, pp. 062006. Cited by: §I.
- [37] NaI(Tl) properties. Note: Online, accessed: 2023-03-30 External Links: Link Cited by: §I.
- [38] (2017) Measurement of the absolute reflectance of polytetrafluoroethylene (PTFE) immersed in liquid xenon. JINST 12, pp. P01017. Cited by: Appendix A.
- [39] (2006) Gamma Ray Spectroscopy with Scintillation Light in Liquid Xenon. JINST 1, pp. P09004. Cited by: §I, §VI.
- [40] (2023) Energy resolution of the LZ detector for high-energy electronic recoils. JINST 8, pp. C04007. Cited by: §I.
- [41] (2011) New measurement of the scintillation efficiency of low-energy nuclear recoils in liquid xenon. Phys. Rev. C 84, pp. 045805. Cited by: §I, §V.
- [42] (2019) Processing of Compton events in the PETALO readout system. In 2019 IEEE Nuclear Science Symposium (NSS) and Medical Imaging Conference (MIC), pp. 1–7. Cited by: §I.
- [43] (2022-05) Monte carlo characterization of PETALO, a full-body liquid xenon-based PET detector. Journal of Instrumentation 17, pp. P05044. Cited by: §I.
- [44] (2025) Evaluation of coincidence time resolution in a liquid xenon detector with silicon photomultipliers. Phys. Rev. Res. 7, pp. 033089. Cited by: §I.
- [45] (2010) Reflectance of polytetrafluoroethylene for xenon scintillation light.. J. Appl. Phys 107 (6), pp. 064902. Cited by: Appendix A.
- [46] (2004) Measurement of the refractive index and attenuation length of liquid xenon for its scintillation light. Nucl. Instrum. Meth. A 516, pp. 462. Cited by: Appendix A, §III.
- [47] (2018) Noble element simulation technique v2.0 (version v2.0.0). Zenodo. Cited by: §I.
- [48] (2020) State of the art in total body PET. EJNMMI Physics 7 (I), pp. 35. Cited by: §I.
- [49] (2014) Advantages of Improved Time Resolution for TOF PET at Very Low Statistics. IEEE Trans. Nucl. Sci. 61 (1), pp. 126 – 133. Cited by: §I.
- [50] (2004) Scintillation response of liquid xe surrounded by ptfe reflector for gamma rays. Nuclear Instruments and Methods in Physics Research Section A: Accelerators, Spectrometers, Detectors and Associated Equipment 535 (3), pp. 692–698. Cited by: §I, §II, §VI.
- [51] (2022) Materials for Future Calorimeters. arXiv. Cited by: §I.